OpenStudy (anonymous):

What is the sum of 9 over the sum of y and 4 + 6 over the difference of y and 3?

OpenStudy (anonymous):

\[9/y+4+6/y+3\]

OpenStudy (anonymous):

\[\frac{\frac{9}{y+4}+6}{y-3}\]

OpenStudy (anonymous):

the 9/y+4 is separate from the 6/y+3

OpenStudy (anonymous):

\[\frac{9}{y+4}+\frac{6}{y-3}\]

OpenStudy (anonymous):

Yes, that's correct

OpenStudy (anonymous):

so wat shud we do||????

OpenStudy (anonymous):

step1; take common denominator

OpenStudy (anonymous):

stay change flip?

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+cb}{bd}\]

OpenStudy (anonymous):

Oh okay so 9(y+3)+6(y+4)/(y+4)(y+3)

OpenStudy (anonymous):

\[{9\over {y+4}}+{6\over{y-3}}={{9(y-3)+6(y+4)}\over{(y+4)(y-3)}}\]

OpenStudy (anonymous):

Okay, then we simplify, correct?

OpenStudy (anonymous):

\[={3(5y-1)}\over{(y+4)(y-6)}\]

OpenStudy (anonymous):

Then you distribute the 3

OpenStudy (anonymous):

and get 15y-3

OpenStudy (anonymous):

haha.. I actually un-distributed it...

OpenStudy (anonymous):

whoops... haha